# paper1.bib

### v1

@@ 5 An exact Quasi-Newton method without exact line search
Assume that the step $\alpha_{k}=1$ is chosen at each iteration of Algorithm 1. Then it is sufficient to build $B_{k}$ using $q_{k}$. There exists a unique value of $\sigma_{r-1}$ such that the algorithm terminates in exactly $r$ iterations. For all other values of $\sigma_{r-1}$ the algorithm will terminate in exactly $r+1$ iterations.
Proof. If $\alpha_{k}=1$ for each iteration of Algorithm 1, then by definition
$p^{N}_{k}$ $=(1-\alpha_{k})p^{N}_{k-1}-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}$
$=-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}//q_{k}.$
Therefore one vector is sufficient to build $B_{k}$.
As shown in Proposition 4.2, there exists a unique scalar $\sigma_{r-1}$ such that $B_{r}$ will generate the Newton step. In this case the algorithm will terminate if unit step is used. Otherwise, $B_{r+1}$ acts as $H$ on the gradient $g_{r+1}$ and therefore the search direction generated will be the Newton step.
Hence, using the correct vectors to create the Hessian approximation $B_{k}$ can lead to termination on (QP) by calculating the Newton step in a finite number of iterations.
@@ 6 Conclusion
In this paper we show sufficient conditions in which to construct a Hessian approximation $B_{k}$ such that exact line search is not necessary to find the Newton step in a finite number of iterations.
The main contribution of this paper is a learning algorithm which can gather information for the Hessian approximation regardless of step sizes that are used as shown in Theorem 5.1. In addition, we show in Corollary 5.1, that without line search, unit step sizes will lead to termination in exactly $r$ or $r+1$ iterations.
Our focus has been to explore what are the requirements for a quasi-Newton method to be exact beyond the usual expectations of conjugate search directions and exact line search. We hope that this can lead to a deeper understanding on the behavior of quasi-Newton methods on general nonlinear problems.

### v2

@@ 5 An exact quasi-Newton method without exact line search
For the unique value $\hat{\sigma}_{r-1}$ stated in Proposition 3.1, Algorithm 1 will generate the Newton step $p_{r-1}=p^{N}_{r-1}(x_{r-1})$, and the algorithm will terminate at iteration $r$ if $\alpha_{r-1}=1$. For all other value of $\sigma_{r-1}$ the algorithm will terminate only when $\alpha_{k}=1$ for some $k\geq r$.
Proof. As stated in Theorem 5.1, for $k\leq r$ the search direction $p_{k}$ coincides with the search direction $p_{k}(x_{k})$ from Proposition 4.1. Therefore, for the unique value of $\hat{\sigma}_{r-1}$ stated in Proposition 3.1
$p_{r-1}=\frac{\hat{\sigma}_{r-1}}{\hat{\sigma}_{r-1}}\hat{q}_{r-1}+p^{N}_{r-2}(x_{r-1})=p^{N}_{r-1}(x_{r-1}).$
Hence the algorithm will terminate at iteration $r$ if $\alpha_{r-1}=1$. For all other values of $\sigma_{r-1}$, the search direction $p_{r-1}$ cannot be parallel to $p^{N}_{r-1}(x_{r-1})$ as $\hat{q}_{r-1}\neq 0$. Therefore the algorithm cannot terminate. Applying Theorem 5.1, for $k\geq r$ the search direction $p_{k}=p^{N}_{r-1}(x_{k})$ is the Newton step, so the algorithm will terminate only if $\alpha_{k}=1$.
[…]
This method will generate the Newton step in a finite number of iterations regardless of step sizes chosen. The action of $B_{k}$ on the correct subspace of two vectors is the vital component which yields the subspace quasi-Newton step of Proposition 4.1. As shown in Lemma A.3 solving $B_{k}p_{k}=-g_{k}$ can be done by solving two systems of linear equations where the matrix is $P_{k}^{T}HP_{k}$, whose dimension is at most $2\times 2$.
@@ 6 Conclusion
In this work we explore exactness in quasi-Newton methods for minimizing a strictly convex quadratic function without the need for the use of exact line search. With knowledge of how the Hessian acts on at most two vectors in the space $\mathcal{K}_{k}(g_{0},H)$, it is possible to generate a quasi-Newton step which is composed of a subspace Newton step to $\hat{x}_{k}$ and a vector parallel to the difference of minimizers $\hat{q}_{k}=\hat{x}_{k+1}-\hat{x}_{k}.$ Using this framework, we have presented an iterative learning algorithm which successively generates such quasi-Newton steps for expanding affine spaces, regardless of step size used.
The algorithm which has been presented generates parallel search directions to that of CG, BFGS and memoryless BFGS when exact line search is used. In the case when exact line search is used and $\sigma_{k}=1$ always, the Hessian approximation $B_{k}$ generated by Algorithm 1 is in fact precisely the approximation that is built using memoryless BFGS. Algorithm 1 can therefore be viewed as a generalization of memoryless BFGS.
The results presented in this work are meant to deepen the theoretical understanding of the action of quasi-Newton methods on quadratic problems. The algorithm presented is a conceptual algorithm to explore the conditions under which it is possible to achieve finite termination on (QP) assuming exact arithmetic. We also hope that our work can lead further research to get a deeper understanding of the behavior of quasi-Newton methods on general nonlinear problems.
Appendix A Appendices
A.1 A first order formulation of Algorithm 1
The construction of $B_{k+1}$ in Algorithm 1 requires information on how $H$ acts on the vectors $q_{k}$ and $p^{N}_{k}$. When a nonzero step is taken at each iteration, this information can be gathered using a difference of gradients. Explicit knowledge of the Hessian $H$ is unnecessary.
Lemma A.1
For $k<r$, if $\alpha_{k}\neq 0$ is chosen in Algorithm 1, then the action of $H$ on $q_{k}$ can be calculated by
$Hq_{k}=\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}$ (A.1)
where $Hp^{N}_{k-1}=B_{k}p^{N}_{k-1}.$ Using this formulation, $p^{N}_{k}$ and $Hp^{N}_{k}$ can be recovered by
$p^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)q_{k}+(1-\alpha_{k})p_{k}$ (A.2)
$Hp^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)Hq_{k}+\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.3)
When $k\geq r$, $p^{N}_{k}=(1-\alpha_{k})p_{k}$, so
$Hp^{N}_{k}=\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.4)
Proof. The step generated by Algorithm 1 is of the form $p_{k}=q_{k}+p^{N}_{k-1}$. Assuming a nonzero step size $\alpha_{k}$ is used so that $x_{k+1}=x_{k}+\alpha_{k}p_{k}$, evaluating the gradient $g_{k+1}$ yields
$\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})=Hp_{k}=Hq_{k}+Hp^{N}_{k-1}.$
Given that the matrix $B_{k}$ acts as $H$ on $p^{N}_{k-1}$, the formulation of $Hq_{k}$ in (A.1) follows directly.
As $p^{N}_{k-1}\in\mathcal{K}_{k}(g_{0},H)$, it is conjugate to $q_{k}$ with respect to $H$. Therefore we may replace the term $q_{k}^{T}Hq_{k}$ by
$q_{k}^{T}Hq_{k}$ $=q_{k}^{T}\left(\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}\right)$
$=\frac{1}{\alpha_{k}}q_{k}^{T}(g_{k+1}-g_{k})$
in (5.1) to obtain the expression (A.2). The formulations for $Hp^{N}_{k}$ in (A.3) and (A.4) follow immediately.
Using the expressions for $p^{N}_{k},Hq_{k}$ and $Hp^{N}_{k}$ given by the lemma above means that Algorithm 1 can be implemented using first order information only, with the same termination guarantees.
A.2 Some linear algebra results
In this appendix, a couple of linear algebra results are given. The first lemma shows that the quasi-Newton matrices generated are positive definite.
Lemma A.2
Let
$B=\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H,$
where $H$ is a symmetric positive definite $n\times n$ matrix, $P$ is an $n\times k$ matrix of rank $k$, and $\sigma$ is a positive constant. Then, $B$ is positive definite.
Proof. We have $B=B_{1}+B_{2}$, for
$B_{1}=\sigma(I-P(P^{T}P)^{-1}P^{T})\text{and}B_{2}=HP(P^{T}HP)^{-1}P^{T}\!H.$
@@ A.2 Some linear algebra results
The following lemma shows a way to solve with the low rank quasi-Newton matrix.
Lemma A.3
Let
$B=\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H,$
where $H$ is a symmetric positive definite $n\times n$ matrix, $P$ is an $n\times k$ matrix of rank $k$, and $\sigma$ is a positive constant. Then, $Bp=-g$ if and only if
$P^{T}\!HP\beta$ $=$ $-P^{T}\!g,$
$P^{T}\!HP\delta$ $=$ $-P^{T}\!Hg-(\sigma P^{T}\!HP+P^{T}\!H^{2}P)\beta,$
$p$ $=$ $-\frac{1}{\sigma}\left(g+P\delta+HP\beta\right).$
Proof. Lemma A.2 shows that $B$ is positive definite, so that $p$ is well defined and unique. In addition, we have
$(\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H)p=-g$
if and only if
$\left(\!\begin{array}[]{ccc}\sigma I&P&HP\\ P^{T}&\frac{1}{\sigma}P^{T}\!P&0\\ P^{T}\!H&0&-P^{T}\!HP\end{array}\!\right)\left(\!\begin{array}[]{c}p\\ \delta\\ \beta\end{array}\!\right)=-\left(\!\begin{array}[]{c}g\\ 0\\ 0\end{array}\!\right)$
for some $k$-dimensional vectors $\delta$ and $\beta$. Elimination of $p$ in the second and third block of equations gives
$\left(\!\begin{array}[]{ccc}\sigma I&P&HP\\ 0&0&-\frac{1}{\sigma}P^{T}\!HP\\ 0&-\frac{1}{\sigma}P^{T}\!HP&-P^{T}\!HP-\frac{1}{\sigma}P^{T}\!H^{2}P\end{array}\!\right)\left(\!\begin{array}[]{c}p\\ \delta\\ \beta\end{array}\!\right)=\left(\!\begin{array}[]{c}-g\\ \frac{1}{\sigma}P^{T}g\\ \frac{1}{\sigma}P^{T}\!Hg\end{array}\!\right).$
We may therefore solve in turn
$P^{T}\!HP\beta$ $=$ $-P^{T}\!g,$
$P^{T}\!HP\delta$ $=$ $-P^{T}\!Hg-(\sigma P^{T}\!HP+P^{T}\!H^{2}P)\beta,$
$p$ $=$ $-\frac{1}{\sigma}\left(g+P\delta+HP\beta\right),$
completing the proof.
This means solving with $B$ can be done by solving two systems of dimension equal to the number of columns in $P$, in our low-rank setting two by two.
[…]